# TWIN-GPT: Digital Twins for Clinical Trials via Large Language Model

### v1

@@ 3.3 Prompt Tasks
This process is repeated each time the patient’s EHR data is updated. The updated EHR data is provided, along with five other EHR profiles similar to the patient’s symptoms, and then adverse events are predicted and medication is recommended again. The goal of this approach is to track and update the patient’s EHR digital twin in real-time, in order to provide more accurate treatment recommendations.
@@
Figure 1: The workflow of TWIN-GPT. (Bottom) TWIN-GPT takes real follow-up visits $X_{n,1:T_{n}}$ of a patient and generates twin visits, $\hat{X}_{n,1:T_{n}}$. (Top) TWIN-GPT uses digital twin $\hat{x}_{n,t}$ to predict the events that occurred in the next timestamp $\hat{x}^{\text{event}}_{n,t+1}$ using a treatment prediction related prompt. We use K nearest neighboring visits for fine-tuning but only use origin visits in prediction. “Treat”: treatment; “Med”: medication; “AE”: adverse event.
3.4 Digital Twin Modeling
To enable TWIN-GPT to learn personalized clinical trial data for patients and to realize data association between different patients, we adopted the method of prompt fine-tuning. In this way, TWIN-GPT is able to accurately digitally twin the characteristics of the patient, and finally realize the prediction of different clinical trials.
Firstly, we fine-tune TWIN-GPT by letting it make predictions based on the patient’s historical access logic so that TWIN-GPT can learn the patient’s access pattern. That is, by predicting the drug treatment and treatment plan at the current time point $t$, the Adverse Event (AE) at the next time point $t+1$ can be predicted so that TWIN-GPT can improve the prediction accuracy of future adverse events by learning the access pattern and drug treatment information in historical data. In particular, in the prediction of AE, we selected $n$ nearest neighbors by the $k$-NN method and processed them accordingly.
At the same time, to realize the association of TWIN-GPT with patient data at different access time points, we fine-tune it to learn the relationship between the data at different time points, better capture the dynamic changes of patient data, and improve TWIN-GPT’s comprehensive and comprehensive understanding of individual patient history. Specifically, we assume the twin process is reversible. That is, there is a mapping relationship that allows you to map the source data space to the twin data space and the twin data space to the source data space.
For each visit $x_{n,t}$, we search all the other patients we retrieve $K$ visits $\left\{x_{k,t^{\prime}}\right\}^{K}_{k}$ from all the other patients with the highest similarity to $x_{n,t}$. Note that $t^{\prime}$ can be either equal or unequal to $t$. Then, we calculated a $x^{\prime}_{n,t}$ based on their similarity, where each number represents the probability of the event occurring. We set $x^{\prime}_{n,t}$ as the input and the $x_{n,t}$ as the target for TWIN-GPT in the twin fine-tuning process. We also fine-tuned the next visit prediction using each patient’s next visit event list, i.e., for the target patient $n$ at visit $t$, using his medication events $x^{\text{med}}_{n,t}$ and treatment events $x^{tr}_{n,t}$ as input and his next adverse events $x^{\text{ae}}_{n,t+1}$ as label for TWIN-GPT in the prediction fine-tuning process. The two fine-tuning processes are carried out alternately so that TWIN-GPT can learn the relationship of all kinds of events.
3.5 Digital Twins Generation
@@ 3.5.1 Personalized Generation
Personalized Generation refers to the process of using EHR data to guide TWIN-GPT in replicating patients, ultimately generating digital twins. Throughout each step of this process, the model incorporates both its data and data from the k-nearest neighbors, enabling TWIN-GPT to engage in in-context learning.
The workflow in Fig. 1 can be summarized in the following key steps: Firstly, we acquire the patient’s self-data (visit) from their EHRs. Next, we perform calculations to determine the $k$-nearest neighbors for this visit, which includes important components such as MED and AE. Subsequently, we apply a masking operation to the self-visit data and design specific prompts for TWIN-GPT. These prompts have two primary tasks: reconstructing the masked data portion and predicting the current treatment strategy.
TWIN-GPT takes over the process, performing both data reconstruction and treatment prediction tasks, ultimately yielding a virtual entity known as “visit:n”. Following this, we provide TWIN-GPT with ground truth information regarding the treatment, enabling it to perform self-corrections and make predictions for the next visit ($visit_{n}+1$).
@@ 3.5.2 Counterfactual Generation
The purpose of this task is to simulate the trajectories of patients under alternative treatment schedules, including switching a patient from the treatment arm (T) to the control arm (C). This simulation not only enhances patient records but also enables the estimation of personalized treatment effects. Moreover, it facilitates the balancing of trial data for predictive modeling, while substantially reducing the necessary sample size for recruiting participants in the control arm. To generate the counterfactual digital twin $\hat{X}_{n,1:T_{n}}$, corresponding to the real patient record $X_{n,1:T_{n}}$, we need to identify the most closely matching patient record within $\tilde{X}_{k,1:T_{k}}$, where $n$ is among $T$ and $k$belongs to $C$. This step merges the unique personal traits of patient $n$ with the temporal patterns observed in patient $k$, facilitating the creation of the synthetic pathway. We evaluate the applicability of TWIN-GPT trajectory generation by comparing the similarity between the synthesized trajectory and the most similar patient record.
4 EVALUATION METRICS
We evaluate the generated data of digital twins based on EHR data from two aspects, aiming at clinical trial outcome prediction and generative privacy.
@@ 4.1 Clinical Trial Outcome Prediction Evaluation of Digital Twins
To assess the practicality of models, we observe whether the additional synthetic data generated by models can successfully predict outcomes under different treatment modalities. Specifically, our experiments involve the following prediction tasks:
@@
To evaluate the similarity between models’ twin data and the real data, we perform a dimensional-wise probability calculation. The calculation of dimension probabilities helps us to understand the probability of occurrence of each feature in the dataset, and it can provide quantitative metrics about the consistency and similarity between synthetic and real data. The Dimension-wise probability (DP) is calculated by
@@
where $V_{f}$ is the number of visits in the dataset containing a specific feature. $V_{l}$ is the total number of accesses in the dataset. When the model successfully produces synthetic data that closely mirrors the original, the distributional properties (DPs) of the synthetic data will approximate those of the real data. We calculate the Person correlation coefficient, $r\in[0,1]$. If $r=1$, it means that the model has high fidelity.
@@ 4.1.3 Severe outcome prediction
To validate whether TWIN-GPT can fully comprehend and replicate the underlying relationship between real-world data records and severe outcomes, we conducted severe outcome prediction validation. Firstly, we defined severe outcomes to include deaths and other critical clinical events. Then, we employed Long Short-Term Memory (LSTM) to forecast these severe outcomes. LSTM received real clinical data and synthetic data generated by different models for model training and prediction. Ultimately, we utilized the Area Under the Receiver Operating Characteristic Curve (AUROC) to assess the accuracy and performance of the predictions.
@@ 4.1.4 Adverse event prediction
Adverse event prediction is to see how much causal relation the synthetic data generated by TWIN-GPT have maintained compared to the raw clinical data. We train another MLP to evaluate the prediction performance, by inputting synthetic data and real data separately.
@@ 4.2.2 Attribute Disclosure
Privacy leakage is when an attacker is able to obtain or infer sensitive information or unknown attributes of an individual, thereby violating an individual’s right to privacy. Attribute leakage occurs when the attacker is able to infer other attributes of the patient based on a known subset of the data. To quantify this risk, we adopt average sensitivity as a metric. Average sensitivity measures the attacker’s ability to infer unknown attributes. It is calculated using the following formula:
@@
where $F_{d}$ represents the number of unknown features discovered, and $F_{l}$ represents the total number of unknown features. For example, if an attacker knows 10 attributes of a patient and can infer 2 unknown attributes from the synthetic data, the average sensitivity would be $2/10=0.2$. This implies that the attacker has a 20% probability of confirming the presence of these unknown attributes in the synthetic data. By computing the average sensitivity, we gain insight into the attacker’s average capability to infer the unknown attributes of the patient. A higher average sensitivity indicates a greater risk of privacy leakage, as it suggests that the attacker has a higher likelihood of inferring unknown attributes. Conversely, a lower average sensitivity indicates a lower risk of privacy leakage, as it implies that the attacker’s inference capability is relatively weak.
@@ 4.2.3 Nearest Neighbor Adversarial Accuracy Risk (NNAA)
Here, $d_{ES}(i)$, $d_{TS}$, $d_{SE}$ and $d_{EE}$, $d_{TT}$, $d_{SS}$ represent the nearest neighbor distances among synthetic data, evaluation data, and real data respectively. For example, if the average distance between synthetic data and evaluation data is greater than that between synthetic data and real data, the NNAA risk score will be positive, indicating potential overfitting of the model to the training data, hence increasing the risk of privacy leakage.
Through these three methods, we can evaluate the privacy risks of synthetic data more comprehensively, thereby better protecting individual privacy.
@@ 5.1.1 Data source
Original clinical trial dataset: We trained TWIN-GPT using a Phase III breast cancer trial dataset (NCT00174655), analyzing Disease-Free Survival (DFS) across different treatment methods. The study randomly divided 2,887 patients into groups to compare the efficacy of docetaxel, either alone or with doxorubicin followed by CMF, against doxorubicin alone or combined with cyclophosphamide, then CMF, in patients with positive axillary lymph nodes. This dataset, publicly available in Project Data, facilitated our performance evaluation. Sphere [^1] https://www.projectdatasphere.org. Trial Outcome Prediction (TOP) dataset: We also utilized the multi-modal TOP dataset [32] for predicting clinical trial outcomes. This dataset integrates clinical trial-related information from multiple data sources, with each trial record including drug molecule information, disease information, eligibility criteria, and trial result information. The dataset is divided into three phases: 1,160 Phase I trials, 4,449 Phase II trials, and 3,436 Phase III trials. The drug molecule information includes the names of candidate drugs and their functional groups. The disease information includes the ICD-10 codes, disease descriptions, and disease hierarchy represented by CCS codes. The eligibility criteria for the trials are described using unstructured natural language, including inclusion and exclusion criteria. The trial result information includes binary indicators for trial success (1) or failure (0), trial phase, start and end dates, trial sponsor, and trial scale (i.e., the number of participants).
@@ 5.1.3 Baseline Models
  • SynTEG [80] is a representative generative model that employs GANs to produce synthetic versions of EHRs. It uses transformer [64] structure in the dependency extraction part and Wasserstein GAN with gradient penalty (WGAN-GP) [3, 38] in the conditional generation part.
@@
  • TWIN-VAE [27] can detailed clinical data through the development of individualized clinical trial digital twins, utilizing variational autoencoders (VAE) for this process.
@@ 5.2 Generation Quality
To make sure that TWIN-GPT has fully learned every dimension’s distribution precisely and accurately, we calculate all DPs of every adverse event separately from both real and generated synthetic data. Fig 3 shows the performance of TWIN-GPT for Breast Cancer data. We compare the performance of dimensions’ distribution of TWIN-GPT and other baselines. The rs of adverse events for EVA, SynTEG, PromptEHR, $k$-NNand TWIN-VAE are -0.06, -0.07, 0.14, 1.0 and 0.99, respectively [27]. We can see that TWIN-GPT has higher fidelity than SynTEG, PromptEHR, and TWIN-VAE. Meanwhile, $k$-NN achieves performance comparable to TWIN-GPT, which can be easily explained that $k$-NN essentially copies and merges fragments from real data to generate synthetic data. However, $k$-NN cannot be used to generate EHR data directly because all the generated data can be linked to their original patient’s record, which will lead to the privacy risk of leaking information. Regarding the TOP dataset, we conducted overall DP calculations for the three medical phases in Fig 3 and independent DPs calculations for the three clinical trial phases as shown in Fig 4. In both scenarios, TWIN-GPT demonstrated high performance with $\geq$0.96. Since the TOP dataset is a multimodal dataset, the comparison algorithms used in the previous dataset do not apply to this dataset. Here, we used $k$-NN to calculate for the TOP dataset, where its $r$ is 1, slightly higher than our TWIN-GPT. As mentioned before, $k$-NN is not suitable due to the potential for privacy risks associated with linking generated data back to the original patient records.
@@
Figure 3: On the TOP dataset, we analyzed the dimension-wise Pearson correlation coefficient (r) of adverse events to evaluate the performance of TWIN-GPT. The x-axis displays the probability across dimensions for real data, while the y-axis signifies the probability associated with synthetic data.
@@
Just like the way we determine the feature-wise fidelity in the Personalized Generation part, we assess the distributional properties (DPs) of actual data from the closest neighbor against the DPs of the corresponding synthetic data to determine the Pearson correlation coefficient ($r$). This comparison yields coefficients that are represented in Fig 6. The results indicate that the digital twins of most patients exhibit a high degree of similarity to their real-life counterparts. These digital twins exhibit high $r$s that are larger than 0.8.

### v2

@@ 3.2 Prompt Tuning for Digital Twin in LLMs
Steps 1-3 are repeated each time the patient’s EHR data is updated. When the updated EHR data is provided, along with another five most similar EHR profiles, the adverse events are predicted and medication is recommended again, as illustrated in table 1 below. The goal of this approach is to track and update the patient’s EHR digital twin in real time, obtaining prompts that most easily establish knowledge associations between input and output.
Visit Input Prompt
$t$
Target visit: $X_{n;1:t-1}$
Five nearest neighbor EHR
Based on the historical visit data of the target patient: xxxxxx, please use the patient’s history and the k most similar visits to generate a digital twin of the target patient and predict the digital twin’s adverse events.
$t+1$
Target visit: $X_{n;1:t}$ EHR
Five nearest neighbor EHR
Based on the historical visit data of the target patient: xxxxxx, please use the patient’s history and the k most similar visits to generate a digital twin of the target patient and predict the digital twin’s recommended medications.
Table 1: Illustration of Prompt Tuning for Digital Twin.
3.3 Digital Twin Generation
In clinical trials, the sequence of real patient visit records is defined as $X_{n;1:T_{n}}$. Our goal is to generate a digital twin $\hat{X}_{n;1:T_{n}}$ that can retain the features of the target patient. For this purpose, we propose a generator, namely TWIN-GPT, that simulates personalized patient trajectories based on the real record $X_{n;1:T_{n}}$. The TWIN-GPT predicts in an auto-regressive manner,
$\hat{X}_{n;T_{n}-1:T_{n}}=\texttt{TWIN-GPT}(X^{\prime}_{n;1:T_{n}-1},\{X_{k;1:T_{k}-1}\}_{k}).$ (5)
In training, $X^{\prime}_{n;1:T_{n}-1}$ is $X_{n;1:T_{n}-1}$, while in test $X^{\prime}_{n;1:T_{n}-1}$ is the previous prediction outcomes $\hat{X}_{n;1:T_{n}-1}$. To enable TWIN-GPT to learn personalized clinical trial data for patients and to leverage data association between different patients, we adopted the method of prompt tuning introduced above. In this way, we hope TWIN-GPT can accurately generate digital twin characteristics of the patient and finally realize the prediction of different clinical trials. We fine-tune TWIN-GPT following the step described in Sec. 3.2, letting it make predictions based on the patient’s historical visits so that TWIN-GPT can learn the patient’s unique characteristics. Specifically, based on the medication ($Med$) and treatment before and planned at the current time point $t$, the Adverse Event ($AE$) at the next time point $t+1$ is predicted. Then based on the $AE_{t+1}$, the $Med_{t+1}$ can be further predicted. Here, the patient’s unique characteristics and information from $k$ similar patients are used as references, to ensure the reliability of the digital twins. Please note that in clinical trials, AEs and medication information are crucial factors we need to focus on, as they significantly impact the success and conclusions of the clinical study. Therefore, our TWIN-GPT is primarily designed to predict these two events.
Notably, different patients may have different trajectories in the trial. To enable TWIN-GPT to create digital twins for various patient patterns, the $k$ most similar patients are retrieved based on all of their previous event sequential $X_{1:t}$. This approach not only helps improve TWIN-GPT’s comprehensive understanding of patient patterns but also prevents it from overfitting to specific individuals. Specifically, at the time step of the visit $x_{n,t}$, we enumerate all the other patients and retrieve $K$ visits $\left\{x_{k,t_{k}}\right\}^{K}_{k=1}$ with the highest cosine similarity to $x_{n,t}$. Note that $t_{k}$ can be either equal or unequal to $t$. By integrating these similar visits as inputs, TWIN-GPT can also enhance its ability to differentiate between similar visits in prediction. The overall workflow is illustrated in Fig. 1.
@@ 3.3 Digital Twin Generation
Figure 1: The workflow of TWIN-GPT. (Bottom) TWIN-GPT takes real follow-up visits $X_{n,1:T_{n}-1}$ of a patient and generates twin visits of next step, $\hat{X}_{n,T_{n}}$. Finally, the whole visit sequence can be predicted. (Top) The top part elaborates on how to use the digital twin visits $\hat{x}_{n,1:t}$ (at the time step $t$) to predict the events that occurred in the next timestamp $\hat{x}^{\text{event}}_{n,t+1}$. We use K nearest neighboring patient visits in TWIN-GPT fine-tuning but only use origin visits in prediction. “Treat”: treatment; “Med”: medication; “AE”: adverse event.
4 APPLICATION & EVALUATION
The generated digital twins are beneficial for simulating clinical trials in silico. In this section, we present two application scenarios (personalized generation and counterfactual generation) and evaluate the quality of the digital twins based on clinical trial outcome prediction accuracy and privacy protection.
4.1 Digital Twins Generation Application
@@ 4.1.1 Personalized Generation
Personalized Generation refers to the process of using EHR data to guide TWIN-GPT in replicating patients, ultimately generating digital twins. Throughout each step of this process, the model incorporates both its data and data from the $k$-nearest neighbors, enabling TWIN-GPT to perform the in-context learning.
@@ 4.1.2 Counterfactual Generation
The purpose of this task is to simulate the trajectories of patients under alternative treatment schedules, including switching a patient from the treatment arm (T) to the control arm (C). This simulation not only enhances patient records but also enables the estimation of personalized treatment effects. Moreover, it facilitates the balancing of trial data for predictive modeling while substantially reducing the necessary sample size for recruiting participants in the control arm. To generate the counterfactual digital twin $\hat{X}_{n,1:T_{n}}$, corresponding to the real patient record $X_{n,1:T_{n}}$, we need to identify the most closely matching patient record within $\tilde{X}_{k,1:T_{k}}$, where $n$ is among $T$ and $k$ belongs to $C$. This step merges the unique personal traits of patient $n$ with the temporal patterns observed in patient $k$, facilitating the creation of the synthetic pathway. We evaluate the applicability of TWIN-GPT trajectory generation by comparing the similarity between the synthesized trajectory and the most similar patient record.
@@ 4.2 Clinical Trial Outcome Prediction Evaluation of Digital Twins
To assess the practicality of digital twin models, we observe whether the additional synthetic data generated by models can successfully predict outcomes under different treatment modalities. Specifically, our experiments involve the following prediction tasks:
@@
To evaluate the similarity between twin data and the real data, we perform a dimensional-wise probability calculation. The calculation of dimension probabilities helps us to understand the probability of occurrence of each feature in the dataset, and it can provide quantitative metrics about the consistency and similarity between synthetic and real data. The Dimension-wise probability (DP) is calculated by
@@
where $V_{f}$ is the number of visits in the dataset containing a specific feature. $V_{l}$ is the total number of accesses in the dataset. When digital twin models successfully produce synthetic data that closely mirrors the original, the distributional properties (DPs) of the synthetic data will approximate those of the real data. We calculate the Person correlation coefficient, $r\in[0,1]$. If $r=1$, it means that the model has high fidelity.
@@ 4.2.3 Severe outcome prediction
To validate whether digital twin models can fully comprehend and replicate the underlying relationship between real-world data records and severe outcomes, we conducted severe outcome prediction validation. Firstly, we defined severe outcomes to include deaths and other critical clinical events. Then, we employed Long Short-Term Memory (LSTM) to forecast these severe outcomes. LSTM received real clinical data and synthetic data generated by different models for model training and prediction. Ultimately, we utilized the Area Under the Receiver Operating Characteristic Curve (AUROC) to assess the accuracy and performance of the predictions.
@@ 4.2.4 Adverse event prediction
Adverse event prediction is to show how much causal relation the synthetic data generated by digital twin models have maintained compared to the raw clinical data. We train another MLP (multiple-layer perceptron) to evaluate the prediction performance by inputting synthetic data and real data separately.
@@ 4.3.2 Attribute Disclosure
Privacy leakage is when an attacker is able to obtain or infer sensitive information or unknown attributes of an individual, thereby violating an individual’s right to privacy. Attribute leakage occurs when the attacker is able to infer other attributes of the patient based on a known subset of the data. To quantify this risk, we adopt average sensitivity as a metric. Average sensitivity measures the attacker’s ability to infer unknown attributes. Mean sensitivity(MS) is calculated using the following formula:
@@
where $F_{d}$ represents the number of unknown features discovered, and $F_{l}$ represents the total number of unknown features. For example, if an attacker knows ten attributes of a patient and can infer two unknown attributes from the synthetic data, the average sensitivity would be $2/10=0.2$. This implies that the attacker has a 20% probability of confirming the presence of these unknown attributes in the synthetic data. By computing the average sensitivity, we gain insight into the attacker’s average capability to infer the patient’s unknown attributes. A higher average sensitivity indicates a greater risk of privacy leakage, as it suggests that the attacker is more likely to infer unknown attributes. Conversely, a lower average sensitivity indicates a lower risk of privacy leakage, as it implies that the attacker’s inference capability is relatively weak.
@@ 4.3.3 Nearest Neighbor Adversarial Accuracy Risk (NNAA)
Here, $d_{ES}(i)$, $d_{TS}$, $d_{SE}$ and $d_{EE}$, $d_{TT}$, $d_{SS}$ represent the nearest neighbor distances among synthetic data, evaluation data, and real data respectively. For example, if the average distance between synthetic data and evaluation data is greater than that between synthetic data and real data, the NNAA risk score will be positive, indicating potential over-fitting of the model to the training data, hence increasing the risk of privacy leakage.
Through these three methods, we can more comprehensively evaluate the privacy risks of synthetic data, thereby better protecting individual privacy.
@@ 5.1.1 Data source
Original clinical trial dataset: We trained TWIN-GPT using a Phase III breast cancer trial dataset (NCT00174655), analyzing Disease-Free Survival (DFS) across different treatment methods. The study randomly divided 2,887 patients into groups to compare the efficacy of docetaxel, either alone or with doxorubicin followed by CMF, against doxorubicin alone or combined with cyclophosphamide, then CMF, in patients with positive axillary lymph nodes. This dataset, publicly available in Project Data, facilitated our performance evaluation.
Trial Outcome Prediction (TOP) dataset: We also utilized the multi-modal TOP dataset [35] for predicting clinical trial outcomes. This dataset integrates clinical trial-related information from multiple data sources, with each trial record including drug molecule information, disease information, eligibility criteria, and trial result information. The dataset is divided into three phases: 1,160 Phase I trials, 4,449 Phase II trials, and 3,436 Phase III trials. The drug molecule information includes the names of candidate drugs and their functional groups. The disease information includes the ICD-10 codes, disease descriptions, and disease hierarchy represented by CCS codes. The eligibility criteria for the trials are described using unstructured natural language, including inclusion and exclusion criteria. The trial result information includes binary indicators for trial success (1) or failure (0), trial phase, start and end dates, trial sponsor, and trial scale (i.e., the number of participants).
@@ 5.1.3 Baseline Models
  • SynTEG [82] is a representative generative model that employs GANs to produce synthetic versions of EHRs. It uses transformer [66] as neural architecture in the dependency extraction part and Wasserstein GAN with gradient penalty (WGAN-GP) [3, 40] in the conditional generation part.
@@
  • TWIN-VAE [30] can detail clinical data through the development of individualized clinical trial digital twins, utilizing variational autoencoders (VAE) for this process.
@@ 5.2 Generation Quality
To make sure that TWIN-GPT has fully learned every dimension’s distribution precisely and accurately, we calculate all DPs of every adverse event separately from both real and generated synthetic data. Fig. 3 shows the performance of TWIN-GPT for Breast Cancer data. We compare the performance of dimensions’ distribution of TWIN-GPT and other baselines. The rs of adverse events for EVA, SynTEG, PromptEHR, $k$-NNand TWIN-VAE are -0.06, -0.07, 0.14, 1.0 and 0.99, respectively [30]. We can see that TWIN-GPT has higher fidelity than SynTEG, PromptEHR, and TWIN-VAE. Meanwhile, $k$-NN achieves performance comparable to TWIN-GPT, which can be easily explained that $k$-NN essentially copies and merges fragments from real data to generate synthetic data. However, $k$-NN cannot be used to generate EHR data directly because all the generated data can be linked to their original patient’s record, which will lead to the privacy risk of leaking information. Regarding the TOP dataset, we conducted overall DP calculations for the three medical phases in Fig. 3 and independent DPs calculations for the three clinical trial phases as shown in Fig. 4. In both scenarios, TWIN-GPT demonstrated high performance with $r\geq$0.96. Since the TOP dataset is a multimodal dataset, the comparison algorithms used in the previous dataset do not apply to this dataset. Here, we used $k$-NN to calculate for the TOP dataset, where its $r$ is 1, slightly higher than our TWIN-GPT. As mentioned before, $k$-NN is not suitable due to the potential for privacy risks associated with linking generated data back to the original patient records.
@@
Figure 3: On the TOP dataset, we analyzed the dimension-wise Pearson correlation coefficient (r) of adverse events to evaluate the performance of TWIN-GPT. The x-axis displays the probability across dimensions for real data, while the y-axis is the probability associated with synthetic data.
@@
Just like the way we determine the feature-wise fidelity in the Personalized Generation part, we assess the distributional properties (DPs) of actual data from the closest neighbor against the DPs of the corresponding synthetic data to determine the Pearson correlation coefficient ($r$). This comparison yields coefficients that are represented in Fig. 6. The results indicate that the digital twins of most patients exhibit a high degree of similarity to their real-life counterparts. These digital twins exhibit high $r$s that are larger than 0.8.